OpenStudy (anonymous):

rational expression x^2-36/6-x

OpenStudy (anonymous):

aroub ?

OpenStudy (anonymous):



OpenStudy (anonymous):

like that ^^^ ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

use: a^2-b^2=(a+b)(a-b) on the top of the fraction

OpenStudy (anonymous):

x^2 -36 = (x)^2 -(6)^2 = (x-6) (x +6) 6-x = - (x-6) (x^2 -36) / (6-x ) = (x-6) (x +6) / -(x-6) = -(x+6) = -x -6

OpenStudy (anonymous):

a=x b=6

OpenStudy (anonymous):